OpenStudy (anonymous):

Estimate 5 1/9-3 5/6

OpenStudy (anonymous):

I got 23/18

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

I need to see how it is work out?

OpenStudy (lucaz):

how did you got 23/18?

OpenStudy (mertsj):

Where are you getting these answers?

OpenStudy (anonymous):

My brother is helping me?

OpenStudy (mertsj):

Did you study the other one where I posted the steps and showed you about the common denominator?

OpenStudy (anonymous):

yes I use it for the 7-3/7 problem

OpenStudy (anonymous):

7/1-3/7=49/7-3/7=46/7 as the answer

OpenStudy (mertsj):

This one is the same thing. Use the common denominator 18

OpenStudy (anonymous):

i still don't know how to work this problem?

OpenStudy (mertsj):

\[\frac{1}{9}= \frac{ }{18}\]

OpenStudy (mertsj):

What is the missing number?

OpenStudy (anonymous):

23

OpenStudy (mertsj):

\[\frac{1}{9}=\frac{23}{18}\]

OpenStudy (mertsj):

Does that look right to you?

OpenStudy (anonymous):

it looks right but how did we get 23

OpenStudy (mertsj):

You really think that 1/9 is the same as 23/18?

OpenStudy (mertsj):



OpenStudy (anonymous):

2*9=18

OpenStudy (mertsj):

What would I do to 9 to change it to 18?

OpenStudy (mertsj):

So if I multiply the denominator by 2, what should I do to the numerator?